OpenStudy (anonymous):

Simplify square root of 3 multiplied by the fifth root of 3 help please

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

\[\sqrt{3} \sqrt[5]{3}=3^{\frac{1}{2}+\frac{1}{5}}=3^{7/10} \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx @robtobey

OpenStudy (anonymous):

Your welcome.